OpenStudy (anonymous):

Solve the equation |k+7|=3

OpenStudy (muzzack):

okay well now take off the absolute value now that it is already isolated and treat it like if it was a normal equation. try it :)

OpenStudy (paki):

k+7= 3 and k+7=-3 solve now

OpenStudy (muzzack):

just like @paki said

OpenStudy (muzzack):

to find out put both the negative and positive forms to see if it is valid/true or false

OpenStudy (anonymous):

OH okay so, if I subtract 7 from there that'll give me -4 right?

OpenStudy (paki):

yup -4.... and what about the other equation...

OpenStudy (anonymous):

That one is -10 i think.

OpenStudy (muzzack):

good :)

OpenStudy (paki):

yeah -10 is rite...

OpenStudy (anonymous):

So the over all answer is (-10,-4) right?

OpenStudy (paki):

you are genius @SparkyTheWolf

OpenStudy (muzzack):

so it is -10 and-4 :)

OpenStudy (muzzack):

i agree @paki

OpenStudy (anonymous):

Okay I understand =-3, but I have others that I need help on also sadly =-(

OpenStudy (paki):

post here...

OpenStudy (muzzack):

go ahead close this question first

OpenStudy (muzzack):

openstudy really doesn't allow many questions in one post, im only enforcing the law but if you dont want to its fine im nice

OpenStudy (paki):

yup agree

OpenStudy (anonymous):

4-3|n+2|>1 Which conjunction or dis junction evaluate the open sentence?

OpenStudy (anonymous):

Oh...I really don't know how...I'm kinda new to this

OpenStudy (muzzack):

hey @paki how are you answer 50 mil questions at a time i just realized right now. SO COOL

OpenStudy (muzzack):

oh easy you know how where it says answer a question and then you posted it underneath it, it says close and bump just click close

OpenStudy (anonymous):

Oh...I'll close it after we solve this one if that is alright.

OpenStudy (muzzack):

3 at the same time

OpenStudy (muzzack):

@paki

OpenStudy (anonymous):

um, can you help me with my other question?

OpenStudy (muzzack):

yeah

OpenStudy (muzzack):

well what is a junction or dis junction do you know what it is

OpenStudy (anonymous):

Disjunction, it made me put a space between the two.

OpenStudy (paki):

4-3|n+2|>1 1 | n+2 | >1 |n+2|>1/1 |n+2|>1 solve now

OpenStudy (muzzack):

well A "logical disjunction" simply means : " or " so; "A ∨ B " is read as "A or B ". Such a disjunction is false if both A and B are false. So, the result would be true if one or more operands are true. It would only be false if both or all it's operands are false. A "logical conjunction" simply means "and" so; the symbol is an inverted "v" " A (inverted v) B " are read as " A and B " . The conjunction would only be true if and only if both operands are true, otherwise it would be false. okay: properties under the conjunctions are: 1. "Associativity": the order of operations does not matter as long as the sequence of the operands is not changed. Eventhough parenthesis are changed, operands will still act the same. ex: (5+1) +5 = 5+ (1+5) 2. "Commutativity": to the ability to change the order of something without changing the end result. ex: 3 + 5 = 5 + 3 for addition; 5*3 = 3*5 for multiplication 3. "Distributivity": uses the "dristibutive law" ex: 5( 2 + 3) = (5*2) + (5*3) 4. "Idempotence": multiple applications of that operation will yield the same result. unary operations: - the absolute value operation is a unary operation on the real numbers - the opposite operation (-x) on the real numbers - the power operations (squaring, cubing, etc) on the real numbers - the factorial operation on the real numbers - the trigonometric functions (sin x, cos x, tan x, cot x, csc x, sec x) on the real numbers - the natural logarithm (ln x) on the real numbers - the logarithm of base 10 (log x) on the real numbers - logical negation on truth values These are examples of unary operations, and could be subject to "Idempotence" 5. "Monotonic" - actually these are used mostly for calculus..

OpenStudy (muzzack):

but @paki just simpled it out hehe all that writing

OpenStudy (anonymous):

|n+2|>1 Okay so I subtract 2 with 1 and got negative -1.

OpenStudy (anonymous):

|n|=-1?

OpenStudy (muzzack):

yes and that means that the absolute value must be bigger than the -2

OpenStudy (muzzack):

i mean -1

OpenStudy (muzzack):

it can't really equal remember it is an inequality

OpenStudy (anonymous):

Okay so it'll look like this |n|>-1

OpenStudy (anonymous):

or n+2>-1

OpenStudy (muzzack):

well actually its better n+2>-1

OpenStudy (anonymous):

Okay, that all I needed help on =-3 thanks